Planet Saturn =5.87 * 10^7 m and it's mass is 5.86 * 10^26 Kg find the density of its material by g/m3
Find it's surface area by m2( 4πr^2)
1
Expert's answer
2019-08-26T10:44:57-0400

The density of the Saturn


"\rho=\frac{M}{V}=\frac{M}{\frac{4}{3}\pi R^3}=\frac{3M}{4\pi R^3}"

"=\frac{3\times 5.86\times 10^{26}\:\rm{kg}}{4\times 3.14\times (5.87 \times 10^7\:\rm{m})^3}"

"=692\:\rm{kg/m^3}=6.92\times 10^5\:\rm{g/m^3}"

The surface area of the Saturn


"A=4\pi R^2"

"=4\times 3.14\times (5.87\times 10^7\:\rm{m})^2=4.33\times 10^{16}\:\rm{m^2}"